# Turnstile area as a measure for chaotic transport in magnetic confinement fusion devices

### v1

@@ Abstract
We analyze stochasticity in the magnetic fields of magnetic confinement fusion reactors by calculating the lobe areas of turnstiles - a method developed for characterizing transport into and out of resonance zones in Hamiltonian dynamical systems. We develop an efficient algorithm based on an action principle to calculate this quantity directly from the magnetic field, including stellarator magnetic fields which are sourced by a complicated set of three-dimensional coils. In the analyzed devices, the turnstile area on the inboard (plasma-facing) manifolds is much smaller than the turnstile area on the outboard (wall-facing) manifolds. The application of the turnstile area calculation for the design of future reactors will be discussed.
@@
Magnetic confinement fusion reactors confine the charged particles of a plasma by channeling them along the twisting field lines of a magnetic field. A magnetic confinement fusion reactor is designed such that in the center, magnetic field lines lie neatly on nested toroidal surfaces; providing good confinement, but at the edge of the plasma the field is designed with features that guide the plasma to specific locations on the wall. The features of the magnetic field that are created at the plasma edge have direct analogies to Hamiltonian Dynamical systems: in a Tokamak a hyperbolic trajectory (x-point) is placed that separates a region of closed trajectories, and guides plasma to the wall along its stable and unstable manifolds. In a stellarator one designs a resonance (magnetic island) at the edge of a region of closed trajectories (plasma region), to control the location where the plasma interfaces with the vessel wall, or one intentionally creates a region of chaotic field lines. In all currently investigated magnetic confinement fusion devices, transport of plasma particles and heat to the wall is partially determined by transport through a homo- or heteroclinic connection at the edge. Transport through such a barrier is quantified by the area of a region of phase space (an amount of magnetic flux) called the ’turnstile’ that under the Hamiltonian flow (field line following) is mapped in or out of the resonance zone (plasma region). In this paper we lift methods to determine this phase-space transport, and apply them to the analysis of the magnetic field in the edge fusion reactors. We demonstrate this method on two-dimensional fast-evaluating maps, on a toy model of a tokamak discharge, and calculate this measure of transport in several stellarator configurations, including configurations of the existing device W7-X. Calculating this transport is important for understanding current and planned fusion reactors, and can be potentially used to control the field structure in the edge as we design future devices.
@@ I Introduction
Magnetic confinement fusion has the goal to generate energy through the fusion isotopes of hydrogen, by heating a plasma confined in a strong magnetic field created by a set of coils. The coils are arranged around a toroidal volume, in which field lines lie on nested surfaces (magnetic surfaces). The charged particles of the plasma are, to lowest order, confined to move along the field lines (integral curves of this field), as their perpendicular motion is constrained through the Lorentz force. Because of this, the effective transport of heat and particles along the field lines is up to ten orders of magnitude higher along the field lines than perpendicular to the field lines, and temperature and pressure are nearly constant on magnetic surfaces.
@@
A good way to understand the particle and heat transport in a fusion reactor is therefore to inspect the trajectories of magnetic field lines. To do this one constructs a Poincaré section, by choosing a cross-section of the torus and recording all the points at which a field line intersects that plane. Such a Poincaré section is shown in in figure 1 for the W7-X stellarator. The trajectories colored in red form the plasma region, where field lines lie on nested magnetic surfaces. At the very center of this foliation is a single field line, called the magnetic axis, named because it is often used as the axis of a toroidal coordinate system. The magnetic axis is a curve that closes on itself after one toroidal turn (long way around the torus). The ratio of the number of times a field line winds around the magnetic axis to the number complete transits of the fusion device is called the rotational transform $\imath$. In tokamaks, which are axisymmetric, it is conventional to use the inverse, called the safety factor $q=1/\imath$.
@@
Figure 1 shows two different configurations that can be achieved in W7-X, the top half shows the standard configuration, the bottom half shows the GYM00+1750 configuration achieved by running different currents through the coils. The trajectories colored in blue form surfaces that do not enclose the magnetic axis, and structures like this are called a magnetic island. These can form on surfaces that have a rational $\imath=m/n$ where $m$ is the number of times field lines on that surface wind around the axis and $n$ the number of full toroidal transits around the device the field line makes to achieve this. In the standard configuration (top half figure 1) the island chain in the edge has $\imath=1$ and consists of 5 islands, and is called a $/5$ 5 island chain. In the GYM00+1750 (bottom panel figure 1) the rotational transform in the edge is higher and $\imath=5/4$. At the center of the islands is a field line that closes on itself, and this is called the island o-point. In between the o-points the island regions approach each other, and there is another field line that closes on itself called the island x-point.
@@
Figure 1: Poincaré section of the magnetic field in the W7-X stellarator. (top): Standard configuration and (bottom): GYM00+1750 configuration with chaotic field lines. Field lines in the core are colored red, field lines in the island divertor are colored blue, and field lines outside of the divertor are colored black and field lines that follow chaotic trajectories are colored grey. The plasma facing components and vessel wall are shown in thick black lines.
@@
W7-X has been designed to have this island chain in the edge, to form what is called an (island) divertor, which creates spatial separation between the plasma facing components and the plasma itself. The wall is also shown in figure 1 with black lines. Since the field lines in the island connect regions close to the plasma with with the outboard side of the island, heat and particles flowing along them can be quickly transported away from the plasma. Conditions along the divertor path (the path the plasma takes from the core to the wall) can be controlled and optimized such that much of the heat is dissipated through recombination and radiation, which minimizes damage to the wall.
@@
Another interesting possibility for the magnetic field topology in the edge of stellarators is the nonresonant divertor [1, 2], in which the central region of nested surfaces is not bounded by a clearly identifyable island chain. Instead the field transitions from nested surfaces to a region of stochastic field lines. Theoretical investigations show that transport through chaotic regions is very important for their operation [3]. Nonresonant divertors are actively being investigated in current experiments  [4] and considered for future devices[5].
@@
There is a vast body of work studying transport in chaotic systems coming from the Hamiltonian dynamical systems’ literature, that can explain a wide-ranging list of physical phenomena. The first step in this regard was the frightening observation by Poincaré that three-body systems such as the earth’s orbit around the sun perturbed by Jupiter are generally chaotic [10]. These fears were laid to rest by theories by Kolmogorov, Arnol’d [11] and Moser [12], proving that invariant orbits persist under small perturbations, providing barriers to the dynamics. The rate of transport through such a barrier when it breaks up can explain the ionization probability of atoms in microwave fields [13]. The probability of passing from one region of phase space to another can also be used to calculate probability of ship capsize [14, 15].
In this paper we focus on a specific quantity that captures the amount of trajectories that pass into and out of a resonance in a dynamical system, called the turnstile (also sometimes called a ’lobe dynamics’). This analysis was developed independently by MacKay, Meiss and Percival [16] and Benmion and Kadanoff [17]. This metric is especially relevant for the magnetic fields in fusion, as the transport into- and out of resonances (the magnetic islands in the edge) directly affects the fusion reactor divertor. We analyze this quantity directly in fusion-relevant fields, and develop an algorithm based on an action principle to efficiently calculate this quantity in the magnetic fields of fusion devices.
@@
This paper is organized as follows: In section II we first describe three types of maps that have been used to describe fusion-relevant fields: iterated maps, perturbed toy models which describe Tokamaks, and the three-dimensional stellarator fields generated by coils‘ In section III we discuss fixed points of the field line map and we describe an algorithm to find them in any map. In section IV we analyze a specific two-dimensional iterated map, calculate the slow- and fast manifolds of several resonances in it, and illustrate how the turnstile mechanism leads to transport in- and out of a resonance zone. We also describe an alogithm to locate hetero- or homoclinic orbits in such maps In section V we develop a toy model describing an axisymmetric tokamak-like magnetic field, and apply perturbations to it. We describe the calculation of the turnstile area based on an action principle, and use it to calculate how the turnstile area changes with perturbation strength. in section VI we apply this to the three-dimensional fields produced by optimized coils of stellarators. We analyze two configurations from the open-source QUASR database, that exhibit very different field structures in the edge. We also analyze a chaotic configuration of the W7-X stellarator called the GYM00+1750 configuration. Section VII concludes with discussion of the competition between chaotic transport in the field and other transport processes, and an outlook for designing future divertors.
@@ II Iterated maps, toy models, and three-dimensional fields
The standard map is an example of a twist map, which equates to fusion magnetic fields where the twist of the field lines of each surface (i.e., the rotational transform $\imath$) is monotonic. 111more precisely, a line constant in an $S^{1}$-valued coordinate is mapped to a line that is single-valued in the other coordinate. Equivalently, a map in which each $m/n$ rational orbit (an orbit that traverses the angle coordinate $n$ times for $m$ applications of the map with $n,m\in\mathbb{N}$, $n\neq 0$) is either part of one continuous family of $m/n$ orbits, only two singular orbits. Aubry [24]-Mather [25] theory applies to twist maps and proves (amongst others) that periodic orbits are related to extrema of an action. Fusion fields however can and sometimes do violate the twist condition, and these can be studied by nontwist maps [26, 6]. These maps also find application in describing particle transport in the fast-particle trajectories [27] (a Hamiltonian system that is different from, though related to the one describing field line flow).
@@ II.2 Perturbed toy models
A next step in complexity comes when we consider fields, and the maps that they generate, that are provided by analytical expressions. There are many different ways to mathematically represent such fields, but a good approach is to start with a formulation that guarantees integrability of the field, and add perturbations to it.
One approach is to exploit the connection to Hamliltonian systems directly, and analytically describe a Hamiltonian. Through a coordinate transformation to a suitable three-dimensional space we get our magnetic field. For example, Punjabi and Boozer have analyzed trajectories Hamiltonian systems that describe a nonresonant divertor [1], and show chaotic mixing of trajectories just around the last closed surface, with several locations where the trajectories quickly leave the system, i.e. are diverted [2, 29]. The locations where the field lines are divertet have been found to be caused by fixed points (see section III[9].
Another method is to describe a simple axisymmetric field, on which we apply perturbations. In axisymmetry (using cylindrical $(R,\phi,Z)$ coordinates) one can construct a flux function $\Psi(R,Z)$, and a separate function describing the toroidal field. The poloidal components of the field (defined as the the components perpendicular to the $\phi$-direction) are defined by treating $\Psi$ as a stream function. Although it is easy to identify the Hamiltonian for such fields, this is not often done. This representation is ubiquitous in tokamak equilibrium reconstruction, where the problem of finding the right flux function can be quickly solved from diagnostic data.
@@ II.3 Maps defined by three-dimensional fields
The methods developed with the iterated maps and the toy model will allow us to analyze chaotic transport in the magnetic fields of actual stellarator fusion reactors. In a stellarator, the magnetic field is generated by a complicated set of coils $\{\mathcal{C}_{i}\}$, and the magnetic field generated by these is calculated from the Biot-Savart law:
@@
where $j_{i}$ is the current carried by coil $\mathcal{C}_{i}$, and $\mathbf{r}$ the point in three dimensions where the field is evaluated, $\mathbf{r}^{\prime}$ is the vector between $\mathbf{r}$ and the point on the coil. Evaluating the Biot-Savart integral is computationally very costly and calculating the field line map can take a significant amount of time per trajectory.
Stellarators are often designed with stellarator symmetry and multiple field periodis $n_{\rm fp}$. This means that there are only a few unique coils that are translated and rotated in space, such that the field repeats itself toroidally with period $T=2\pi/n_{\rm fp}$. This means that the field line map can be evaluated more efficiently by only integrating over a single period:
@@
We note that equation (3) requires that $B^{\phi}<0$ everywere (or $B^{\phi}>0$). The fields defined by a finite set of coils, e.g. equation (2) do not necessarily fulfil this requirement, and in a given section there can exist field lines that do flow around the torus, but for example get caught close to a coil. This means that the field line map cannot be defined at every point of $(R,Z)\in(\mathbb{R}^{+}\times\mathbb{R})$, but we must restrict the field line map to a region [9], called the maximal mappable region.
@@
III The field line map and it’s fixed points and manifolds.
@@
A fixed point of the field line $n$-map is a point where $f^{n}(R,Z)=(R,Z)$. The existence of at least one fixed points is guaranteed by Brouwers’ fixed point theorem [32], assuming that there is at least some region that maps wholly onto itself. Given that the goal of a magnetic confinement fusion reactor is to generate a field where field lines foliate nested toroidal surfaces (invariant tori), this is a necessary condition for the fields under consideration. The ‘magnetic axis’, the field line that lies at the center of this foliation of magnetic surfaces, is a fixed point of the field line map.
@@
Fixed points can be located using a Newton’s method. To first order, using the linearized map around a point $\mathbf{x}$ the field line map can be approximated as:
$f^{n}(\mathbf{x}+\delta\mathbf{x})\approx f^{n}(\mathbf{x})+\mathsf{M}\delta\mathbf{x}$ (4)
where the matrix $\mathsf{M}$ (with indices $\mathsf{M}_{ij}=\partial_{j}f^{n}_{i}$, the labels $i,j$ run over the dimensions $R$ and $Z$) is the Jacobian matrix of the field line map. $\mathsf{M}$ is a $2\times 2$ matrix with real coefficients. From this, we can solve for which step $\delta\mathbf{x}$ the linear approximation brings us to a fixed point, i.e. $f(\mathbf{x}+\delta\mathbf{x})=\mathbf{x}+\delta\mathbf{x}$:
$\mathbf{x}+\delta\mathbf{x}$ $=f^{n}(\mathbf{x})+\mathsf{M}\delta\mathbf{x}\implies$ (5)
$\mathsf{M}\delta\mathbf{x}-\mathbb{I}\delta\mathbf{x}$ $=\mathbf{x}-f(\mathbf{x}).$ (6)
@@
$\mathbf{x}_{k+1}=\mathbf{x}_{k}+(\mathsf{M}-\mathbb{I})^{-1}(\mathbf{x}-f(\mathbf{x}_{k})).$ (7)
@@
$\partial_{j}f^{n}_{i}=\int_{phi_{0}}^{\phi_{0}+nT}\partial_{j}\frac{B^{i}}{B^{\phi}}\mathrm{d}\phi=\int_{phi_{0}}^{\phi_{0}+nT}\frac{\partial_{j}B^{i}}{B^{\phi}}-\frac{B^{i}}{(B^{\phi})^{2}}\partial_{j}B^{\phi}\mathrm{d}\phi$ (8)
starting with initial value $\partial_{j}f^{n=0}_{i}=\mathbb{I}$, and where the trajectory of integration is determined by the field line itself, equation (3).
@@
The set of all trajectories that asymptotically approach a hyperbolic point ($\{\mathbf{x}|\lim_{m\rightarrow\infty}f^{nm}(\mathbf{x})=\mathbf{x}_{0}\}$) are called the stable manifold. In the linear region around the fixed point (where eq. (4) holds), the manifold is the span of $\mathbf{v}_{2}$, as the less-than unity eigenvalue implies that points get mapped closer, but the manifold extends much further as will be discussed in the next section. The set of trajectories that asymptotically leave the fixed point, or conversely, whose limit $\{\mathbf{x}|\lim_{m\rightarrow-\infty}f^{nm}(\mathbf{x})=\mathbf{x}_{0}$ are called the unstable manifold, which starts off in the linear regime as the span of $\mathbf{v}_{1}$.
@@
Figure 2: Trajectories of a Tokamap configuration, showing an inner region of mostly closed surfaces, bordered by a stochastic region in which transport is determined by field structures surrounding large resonances. 80 trajectories are followed for 4000 iterations each. The unstable fixed points in the 1/3, 2/5 and 1/2 island are shown in blue, and their stable manifolds are shown in blue, their unstable manifolds are shown in magenta. The stable (unstable) manifolds associated with these fixed points are shown in magenta (blue)
@@
We have plotted trajectories of the Tokamap in figure 2, with the values $K=4.7$ and $w_{0}=0.9$. In the vicinity of $\psi=0$ there is a region where trajectories close in $\vartheta$, corresponding to closed magnetic surfaces. The trajectories are given a gradient in color starting with yellow at $\psi=0$ and $\vartheta=0.5$ to dark blue at $\psi=0.43$, $\vartheta=0$. At larger $\psi$, we see trajectories that do not close in $\vartheta$, but form several discrete circles. These trajectories are analogous to magnetic islands. The stochasticity in the field increases until a large period 3 island chain is seen, which forms the boundary to a completely stochastic field in which a period 5 island chain and a period 2 island chain are seen.

### v3

@@ Abstract
We analyze a measure for the degree of chaos in the magnetic fields of magnetic confinement fusion reactors by calculating the lobe areas of turnstiles - a method developed for characterizing transport into and out of resonance zones in Hamiltonian dynamical systems. We develop an efficient algorithm based on an action principle to calculate this quantity directly from the magnetic field, including stellarator magnetic fields which are sourced by a complicated set of three-dimensional coils. In the analyzed devices, the turnstile area on the inboard (plasma-facing) manifolds is smaller than the turnstile area on the outboard (wall-facing) manifolds. This indicates that the relative importance of chaos increases as the plasma facing components are moved further away from the plasma. The application of the turnstile area calculation for the design of future reactors will be discussed.
@@
Magnetic confinement fusion reactors confine the charged particles of a plasma by channeling them along the twisting field lines of a magnetic field. A magnetic confinement fusion reactor is designed such that in the center, magnetic field lines lie neatly on nested toroidal surfaces; providing good confinement, but at the edge of the plasma the field is designed to guide the plasma to specific locations on the chamber wall. Transport of plasma particles and heat to the wall is largely determined by the amount of field lines (amount of magnetic flux) connecting the region around the plasma to the region near the wall. A quantity called the ’turnstile area’, which is zero if a field is not chaotic and increases as the chaos increases, has been defined in the study of Hamiltonian dynamical systems to quantify the amount of trajectories that connect one region of phase space to another. In this paper we use the correspondence between magnetic fields and Hamiltonian dynamical systems, and apply this method to the analysis of the magnetic field in the edge of fusion reactors. We demonstrate this method on two-dimensional fast-evaluating maps, on a toy model of a tokamak, and calculate this measure of transport in several stellarator configurations, including configurations of the existing device W7-X. Calculating this transport is important for understanding current and planned fusion reactors, and can be potentially used to control the field structure in the edge as we design future devices.
@@ I Introduction
Magnetic confinement fusion has the goal to generate energy through the fusion isotopes of hydrogen, by heating a plasma confined in a strong magnetic field created by a set of coils. The coils are arranged around a toroidal volume, in which field lines lie on nested surfaces (magnetic surfaces). The charged particles of the plasma are, to lowest order, confined to move along the field lines (integral curves of this field), as their perpendicular motion is constrained through the Lorentz force. Because of this, the effective transport of heat and particles along the field lines is up to ten orders of magnitude higher [1] along the field lines than perpendicular to the field lines, and temperature and pressure are nearly constant on magnetic surfaces.
@@
A good way to understand the particle and heat transport in a fusion reactor is therefore to inspect the trajectories of magnetic field lines. To do this one constructs a Poincaré section, by choosing a cross-section of the torus and recording all the points at which a field line intersects that plane. Such a Poincaré section is shown in in figure 1 for the W7-X stellarator. The trajectories colored in red form the plasma region, where field lines lie on nested magnetic surfaces. At the very center of this foliation is a single field line, called the magnetic axis, named because it is often used as the axis of a toroidal coordinate system. The magnetic axis is a curve that closes on itself after one toroidal turn (long way around the torus). The ratio of the number of times a field line winds around the magnetic axis to the number of complete transits of the fusion device is called the rotational transform $\imath$. In tokamaks, which are axisymmetric, it is conventional to use the inverse, called the safety factor $q=1/\imath$.
@@
Figure 1 shows two different configurations that can be achieved in W7-X, the top half shows the standard configuration, the bottom half shows the GYM00+1750 configuration achieved by running different currents through the coils. The trajectories colored in blue form surfaces that do not enclose the magnetic axis, and structures like this are called a magnetic island. These can form on surfaces that have a rational $\imath=m/n$ where $m$ is the number of times field lines on that surface wind around the axis and $n$ the number of full toroidal transits around the device the field line makes to achieve this. In the standard configuration (top half figure 1) the island chain in the edge has $\imath=1$ and consists of 5 islands, and is thus an $m/n=5/5$ island chain. In the GYM00+1750 (bottom panel figure 1) the rotational transform in the edge is higher and $\imath=5/4$. At the center of the islands is a field line that closes on itself, and this is called the island o-point. In between the o-points the island regions approach each other, and there is another field line that closes on itself called the island x-point.
@@
Figure 1: Poincaré section of the magnetic field in the W7-X stellarator. (top): Standard configuration and (bottom): GYM00+1750 configuration with chaotic field lines. Field lines in the core are colored red, field lines in a magnetic island surrounding the core are colored blue, field lines outside of the divertor are colored black and field lines that follow chaotic trajectories are colored gray. The plasma facing components and vessel wall are shown in thick black lines.
@@
W7-X has been designed to have this island chain in the edge, to form what is called an (island) divertor, which creates spatial separation between the plasma facing components and the plasma itself. The plasma facing components (PFCs) are also shown in figure 1 with black lines. Since the field lines in the island connect regions close to the plasma with the outboard side of the island, heat and particles flowing along them can be quickly transported away from the plasma. Conditions along the divertor path (the path the plasma takes from the core to the PFCs) can be controlled and optimized such that much of the power is dissipated through recombination and radiation, which minimizes damage to the PFCs [2].
When the pressure in the confined plasma is increased, the magnetic field changes, and the field lines around the x-points do not lie on neat surfaces, but become chaotic [3]. There are other configurations in W7-X, generated by running different coil currents, that have chaotic field lines in the edge even when there is no pressure. One of these is the GYM00-1750 configuration that will be analyzed in this paper. A first step towards understanding the transport of heat and particles through such chaotic regions is to calculate the amount of field lines that connect the region close to the plasma to the regions near the PFCs.
@@
There has also been discussion in the literature about the concept of a nonresonant divertor [4, 5], in which the central region of nested surfaces is not bounded by a clearly identifiable island chain. Instead the field transitions from nested surfaces to a region of chaotic field lines. Theoretical investigations show that transport through chaotic regions is very important for their operation [6]. Nonresonant divertors are actively being investigated in current experiments  [7] and considered for future devices[8].
In a tokamak fusion reactor, the main coils that confine the plasma are axisymmetric. Axisymmetry precludes the formation of chaotic trajectories, but error fields due to coil misalignment [9] can introduce some chaos. Other tokamak experiments intentionally add non-axisymmetric coils. In TEXTOR-DED coils were added to produce an ‘ergodic divertor‘ [10]. The tokamaks Tore-Supra [11], DIIID[12], KSTAR [13] and ITER [14] (among others) have nonaxisymmetric ‘Resonant Magnetic Perturbation‘ (RMP) [15] coils which create a region of chaotic field at the plasma edge. The transport of heat and particles through the generated chaotic region is actively being studied to predict divertor performance in ITER [16, 17]. RMP fields cause structures called ‘homoclinic tangles’ generated by the field lines which asymptotically limit to the tokamak x-point. These can be explicitly calculated in fusion devices [18], and are seen in experiments [19, 20].
In the case of the GYM+1750 configuration we see a clear example where chaos influences divertor behavior. To reach the PFCs, the plasma must be somehow transported through a chaotic region. How the heat/particles reach the PFCs (i.e. the structure of the transport, and hence the temperature and density distribution in the edge) is not immediately obvious from the Poincaré section. The same occurs in perturbed tokamaks, and in nonresonant divertors. Though other transport processes exist, such as neoclassical transport [21] and turbulent transport [22], parallel transport is likely to be important due to the very high heat conductivity along the field lines. If we want to quantitatively understand and/or optimize for/against chaos, we need a robust and quantitative toolkit of chaos analysis.
In this paper we will analyze transport through a chaotic region in fusion-relevant configurations using methods lifted from Hamiltonian dynamical systems. We will calculate structures similar to [18], and provide an efficient algorithm to calculate the flux enclosed by the structures. We will apply these methods to simple maps, to tokamak-like fields and to stellarator fields.
@@
In this paper we focus on a specific quantity that captures the amount of trajectories that pass into and out of a resonance in a dynamical system, called the turnstile (also sometimes called a ’lobe dynamics’). This analysis was developed independently by MacKay, Meiss and Percival [33] and Benimon and Kadanoff [34]. This metric is especially relevant for the magnetic fields in fusion, as the transport into- and out of resonances (the magnetic islands in the edge) directly affects the fusion reactor divertor. We analyze this quantity directly in fusion-relevant fields, and develop an algorithm based on an action principle to efficiently calculate this quantity in the magnetic fields of fusion devices.
@@
This paper is organized as follows: In section II we first describe three types of maps that have been used to describe fusion-relevant fields: iterated maps, toy models which describe perturbed tokamaks, and the three-dimensional stellarator fields generated by coils. In section III we discuss fixed points of the field line map and we describe an algorithm to find them in any map. In section IV we analyze a specific two-dimensional iterated map, calculate the stable- and unstable manifolds of several resonances in it, and illustrate how the turnstile mechanism leads to transport in- and out of a resonance zone. We also describe an algorithm to locate hetero- or homoclinic orbits in such maps. In section V we develop a toy model describing an axisymmetric tokamak-like magnetic field, and apply perturbations to it. We describe the calculation of the turnstile area based on an action principle, and use it to calculate how the turnstile area changes with perturbation strength. In section VI we apply this to the three-dimensional fields produced by optimized coils of stellarators. We analyze two configurations from the open-source QUASR database [35, 36], that exhibit very different field structures in the edge. We also analyze a chaotic configuration of the W7-X stellarator called the GYM00+1750 configuration. Section VII concludes with discussion of the competition between chaotic transport in the field and other transport processes, and an outlook for future work.
@@ II Iterated maps, toy models, and three-dimensional fields
The standard map is an example of a twist map, which equates to fusion magnetic fields where the twist of the field lines of each surface (i.e., the rotational transform $\imath$) is monotonic. 111more precisely, a line constant in an $S^{1}$-valued coordinate is mapped to a line that is single-valued in the other coordinate. Equivalently, a map in which each $m/n$ rational orbit (an orbit that traverses the angle coordinate $n$ times for $m$ applications of the map with $n,m\in\mathbb{N}$, $n\neq 0$) is either part of one continuous family of $m/n$ orbits, only two singular orbits. Aubry [43]-Mather [44] theory applies to twist maps and proves (among others) that periodic orbits are related to extrema of an action. Fusion fields however can and sometimes do violate the twist condition, and these can be studied by nontwist maps [45, 23]. These maps also find application in describing particle transport in the fast-particle trajectories [46] (a Hamiltonian system that is different from, though related to the one describing field line flow).
@@ II.2 Perturbed toy models
A next step in complexity comes when we consider magnetic fields, and the maps that they generate, that are provided by analytical expressions. There are many different ways to mathematically represent such fields, but a good approach is to start with a formulation that guarantees integrability of the field, and add perturbations to it.
One approach is to exploit the connection to Hamiltonian systems directly, and analytically describe a Hamiltonian. Through a coordinate transformation to a suitable three-dimensional space we get our magnetic field. For example, Punjabi and Boozer have analyzed trajectories in Hamiltonian systems that describe a nonresonant divertor [4], and show chaotic mixing of trajectories just around the last closed surface, with several locations where the trajectories quickly leave the system, i.e. are diverted [5, 48]. The locations where the field lines are diverted have been found to be caused by fixed points (see section III[26].
Another method is to describe a simple axisymmetric field, on which we apply perturbations. In axisymmetry (using cylindrical $(R,\phi,Z)$ coordinates) one can construct a flux function $\Psi(R,Z)$, and a separate function describing the toroidal field. The poloidal components of the field (defined as the components perpendicular to the $\phi$-direction) are defined by treating $\Psi$ as a stream function. Although it is easy to identify the Hamiltonian for such fields, this is not often done. This representation is ubiquitous in tokamak equilibrium reconstruction, where the problem of finding the right flux function can be quickly solved from diagnostic data.
@@ II.3 Maps defined by three-dimensional fields
In a stellarator, the magnetic field is generated by a complicated set of coils $\{\mathcal{C}_{i}\}$, and the magnetic field generated by these is calculated from the Biot-Savart law:
@@
where $j_{i}$ is the current carried by coil $\mathcal{C}_{i}$, $\mathbf{r}$ the point in three dimensions where the field is evaluated, and $\mathbf{r}^{\prime}$ is the vector between $\mathbf{r}$ and the point on the coil. Evaluating the Biot-Savart integral is computationally very costly and calculating the field line map can take a significant amount of time per trajectory.
Stellarators are often designed with stellarator symmetry [49] and multiple field periods $n_{\rm fp}$. This means that there are only a few unique coils that are translated and rotated in space, such that the field repeats itself toroidally with period $T=2\pi/n_{\rm fp}$. This means that the field line map can be evaluated more efficiently by only integrating over a single period:
@@
We note that equation (3) requires that $B^{\phi}<0$ everywhere (or $B^{\phi}>0$). The fields defined by a finite set of coils, e.g. equation (2) do not necessarily fulfill this requirement, and in a given section there can exist field lines that do flow around the torus, but for example get caught close to a coil. This means that the field line map cannot be defined at every point of $(R,Z)\in(\mathbb{R}^{+}\times\mathbb{R})$, but we must restrict the field line map to a region [26], called the maximal mappable region.
@@
III The field line map and its fixed points and manifolds.
@@
A fixed point of the field line $n$-map is a point where $f^{n}(R,Z)=(R,Z)$. The existence of at least one fixed point is guaranteed by Brouwers’ fixed point theorem [50], assuming that there is at least some region that maps wholly onto itself. Given that the goal of a magnetic confinement fusion reactor is to generate a field where field lines foliate nested toroidal surfaces (invariant tori), this is a necessary condition for the fields under consideration. The ‘magnetic axis’, the field line that lies at the center of this foliation of magnetic surfaces, is a fixed point of the field line map.
@@
Fixed points of the map applied $n$ times can be located using a Newton’s method. To first order, using the linearized map around a point $\mathbf{x}$ the field line map can be approximated as:
$f^{n}(\mathbf{x}+\delta\mathbf{x})\approx f^{n}(\mathbf{x})+\mathsf{M}^{n}\delta\mathbf{x}$ (4)
where the matrix $\mathsf{M}^{n}$ (with indices $\mathsf{M}_{ij}=\partial_{j}f^{n}_{i}$, the labels $i,j$ run over the dimensions $R$ and $Z$) is the Jacobian matrix of the field line map. $\mathsf{M}$ is a $2\times 2$ matrix with real coefficients. From this, we can solve for which step $\delta\mathbf{x}$ the linear approximation brings us to a fixed point, i.e. $f^{n}(\mathbf{x}+\delta\mathbf{x})=\mathbf{x}+\delta\mathbf{x}$:
$\mathbf{x}+\delta\mathbf{x}$ $=f^{n}(\mathbf{x})+\mathsf{M}^{n}\delta\mathbf{x}\implies$ (5)
$\mathsf{M}^{n}\delta\mathbf{x}-\mathbb{I}\delta\mathbf{x}$ $=\mathbf{x}-f(\mathbf{x}).$ (6)
@@
$\mathbf{x}_{k+1}=\mathbf{x}_{k}+(\mathsf{M}^{n}-\mathbb{I})^{-1}(\mathbf{x}-f^{n}(\mathbf{x}_{k})).$ (7)
@@
$\partial_{j}f^{n}_{i}=\int_{\phi_{0}}^{\phi_{0}+nT}\partial_{j}\frac{B^{i}}{B^{\phi}}\mathrm{d}\phi=\int_{\phi_{0}}^{\phi_{0}+nT}\frac{\partial_{j}B^{i}}{B^{\phi}}-\frac{B^{i}}{(B^{\phi})^{2}}\partial_{j}B^{\phi}\mathrm{d}\phi$ (8)
starting with initial value $\partial_{j}f^{n=0}_{i}=\delta_{ij}$ (Kronecker $\delta$), and where the trajectory of integration is determined by the field line itself, equation (3).
@@
The set of all trajectories that asymptotically approach a hyperbolic point ($\{\mathbf{x}|\lim_{m\rightarrow\infty}f^{nm}(\mathbf{x})=\mathbf{x}_{0}\}$) are called the stable manifold. In the linear region around the fixed point (where eq. (4) holds), the manifold is the span of $\mathbf{v}_{2}$, as the less-than unity eigenvalue implies that points get mapped closer, but the manifold extends much further as will be discussed in the next section. The set of trajectories that asymptotically leave the fixed point, or conversely, whose limit $\{\mathbf{x}|\lim_{m\rightarrow-\infty}f^{nm}(\mathbf{x})=\mathbf{x}_{0}$ are called the unstable manifold, which starts off in the linear regime as the span of $\mathbf{v}_{1}$. Since plasma can follow the field lines in both directions, heat and plasma can be transported both towards and away from the hyperbolic fixed point along both manifolds.
@@
Figure 2: Trajectories of a Tokamap configuration, showing an inner region of mostly closed surfaces, bordered by a chaotic region in which transport is determined by field structures surrounding large resonances. 80 trajectories are calculated for 5000 iterations each. The hyperbolic fixed points of the period 3, period 5 and period 2 island are shown in black. The manifolds of these hyperbolic fixed points are shown in blue (for the stable manifolds) and magenta (for the unstable fixed point).
@@
We have plotted trajectories of the Tokamap in figure 2, with the values $K=4.7$ and $w_{0}=0.9$. In the vicinity of $\psi=0$ there is a region where trajectories close in $\vartheta$, corresponding to closed magnetic surfaces. The trajectories are given a gradient in color starting with yellow at $\psi=0$ and $\vartheta=0.5$ to dark blue at $\psi=0.43$, $\vartheta=0$. At larger $\psi$, we see trajectories that do not close in $\vartheta$, but form several discrete circles. These trajectories are analogous to magnetic islands. The amount of chaotic trajectories in the field increases until a large period 3 island chain is seen, which forms the boundary to a very chaotic field in which a period 5 island chain and a period 2 island chain are seen.